OpenStudy (anonymous):

Assume that y varies inversely with x. If y = -3 when x = 2.4, what is y when x = 2?

OpenStudy (nirmalnema):

x1y1=x2y2 2.4*(-3)=2*y y=-7.2/2=-3.6

OpenStudy (anonymous):

Okay, how do you solve a problem like that? Like, whats the formula?

OpenStudy (anonymous):

yx = k -> (-3)(2.4) = k y(2) = (-3)(2.4) -> y = (-3)(2.4)/2

OpenStudy (nirmalnema):

x1=1/y1 x2=1/y2 so x1y1=1 x2y2=1 therefore x1y1=x2y2

OpenStudy (nirmalnema):

@Cubi-Cal have u got it

OpenStudy (anonymous):

Yeah I think so... let me double check.

OpenStudy (nirmalnema):

ok.....

OpenStudy (anonymous):

The formula for what I wrote is as follows: You have an inverse proportion, therefore: y times x will equal a constant. You are given an initial "y" and "x", so you can derive the constant "k".

OpenStudy (anonymous):

Now, "y" times some "x" will give you a known "k" (constant), determined from the first step. You have "x" and "k", so you can determine the new "y".

OpenStudy (anonymous):

All good now, @Cubi-Cal ?

OpenStudy (anonymous):

Yessir! (Ma'am) Thanks you two!:)

OpenStudy (anonymous):

uw! @Cubi-Cal

OpenStudy (anonymous):

Wait, would x=-7.2 or -3.6?

OpenStudy (nirmalnema):

x = -3.6